OpenStudy (anonymous):

sqrt 250 x^16 / sqrt 2x

OpenStudy (kuoministers):

hi :)

OpenStudy (anonymous):

hey! lol can you help me with this problem

OpenStudy (kuoministers):

\[ \frac{ \sqrt{250x ^{16}} }{ \sqrt{2x }}\] This is the question right? meoowwwwww

OpenStudy (anonymous):

yep!

OpenStudy (kuoministers):

so we need to simply this right?

OpenStudy (anonymous):

yep!!

OpenStudy (kuoministers):

what 2 numbers can u break down 250 like using numbers such as 4, 9, 36, 81(theres a lot more)

OpenStudy (kuoministers):

you know how to simply a surd e.g sqrt 12 is 2 sqrt 3

OpenStudy (anonymous):

would it be 10 and 25?

OpenStudy (kuoministers):

:) yes!!

OpenStudy (kuoministers):

and sqrt x^16 would be? e.g sqrt x^14 is 7x

OpenStudy (anonymous):

4x?

OpenStudy (kuoministers):

hmm not really its half.. like sqrt x^2 is only x which is half right? so it would be 8x

OpenStudy (anonymous):

oh alright ya that makes more since!

OpenStudy (anonymous):

so what would the equation look like as of right now?

OpenStudy (kuoministers):

actually the sqrt x^14 was wrong :P

OpenStudy (anonymous):

so what would it be then? lol

OpenStudy (kuoministers):

\[\sqrt{x^{14}} \] would be..

OpenStudy (kuoministers):

\[x \sqrt{x^{7}}\] cause x^2 times x^7 gives you sqrt x^14

OpenStudy (anonymous):

so would \[\sqrt[x]{x^8}\]

OpenStudy (anonymous):

or i mean x sqrt x^8

OpenStudy (kuoministers):

\[\frac{ \ 5x ^{8} { \sqrt{10 ^{}}}{ } }{ \sqrt{2x} }\]

OpenStudy (kuoministers):

wow that took forever :P meoowwww

OpenStudy (anonymous):

lol so thats the final answer?

OpenStudy (kuoministers):

ok my bad but i remembered!!! sqrt x^16 is x^8!!! and sqrt x^14 is x^7!!!

OpenStudy (kuoministers):

it should be!!! if not pls tell me :P but it should be right

OpenStudy (anonymous):

Alright thank you a lot!

OpenStudy (kuoministers):

np!! dont forgot the medal :D

OpenStudy (kuoministers):

and thanks to you for being patient with a cat

OpenStudy (anonymous):

np! and i wont1

OpenStudy (anonymous):

wont! lol